Find the area K of the triangle specified below. a = 10, b= 8, c = 6 The area K is square units. (Type an integer.)
Answer #1

a=10, b = 8, C 6 we know herons formula of triangles area ✓ scs-a7 Cs-b2c5-4) where s = a+b+c 2 So here s = 104 846 12 2 so

Thumbs up please